pls. help me solve this problem.... thank you.
problem 1. The area of the rectangle is 350 sq. cm. If the length exceeds 3 times its width by 5, find the dimensions of the rectangle.
problem 2. If the length of a certain rectangle were invreased by 2 ft and the width by 1 foot, the area would be decreased by 2 square feet. I f the length were decresed by 2 ft, the area would be increased by 16 square fet. Find the dimensions of the rectangle.

Answer by mananth(16946)

let the width be x
the length = 3x+5
area = 350 sq.m
..
area = L*W
x(3x+5)=350
3x^2+5x=350
3x^2+5x-350=0
using quadratic formula -5 (+/-) sqrt b^2-4ac /2a
discriminant = 4225
= -5+sqrt 4225 / 6
=-5+65 /6
= 60/6
= 10 cm
the other root will be negative so ignore the value
length = 3x+5 = 35 cm
...
m.ananth@hotmail.ca

=